can some one please help me Solve the system using addition:
x/2 + y = 5/2
x/3 -y/3 = -1/3
I cannot seem to figure this out I am stumped on how to get opposite coefficients.

Answer by Alan3354(69443)

x/2 + y = 5/2
x/3 -y/3 = -1/3
I cannot seem to figure this out I am stumped on how to get opposite coefficients.
=============
IDK what "opposite coefficients" are.
================
Multiply the 2nd eqn by 3, then add them